$500 is deposited in an account which bears 12%p.a interest compound monthly. Use the compound interest formula to find the number of years it would take to accumulate $500 to $806.11

1
Expert's answer
2022-03-25T15:24:01-0400

"A=P(1+\frac{r}{n})^{n×t}"

for compounded monthly n=12

"806.11=500(1+\frac{0.12}{12})^{12×t}"


"806.11=500(1.01)^{12t}"

"806.11\/500=(1.01)^{12t}"

"1.61222=(1.01)^{12t}"

log1.61222="12tlog1.01"

"12t=log1.61222\/log1.01=48.0"

"t=48\/12=4yrs"